# Leveraging Large Language Models for NLG Evaluation: A Survey

### v1

@@ 4 Benchmarks and Tasks
5 Open Problems
@@
LLMs have been expanded w.r.t their broad capabilities beyond traditional single-task focuses, encompassing complex instructions like coding and open-ended real-world requirements OpenAI (2023); Significant Gravitas . Consequently, there is an increasing demand for more comprehensive and flexible evaluation methods. However, traditional evaluation methods and most current LLM-based evaluators are limited to constrained tasks and evaluation aspects (cf. Table 2). Some promising attempts have been made in this direction. For instance, MT-Bench Zheng et al. (2023) uses GPT-4 as an evaluator across multiple domains for multi-turn questions. Yet, this is too confined to a few evaluation aspects and limits dialogue to two turns only. Another model, Auto-J Li et al. (2023b), approaches from a data construction perspective, training a 13B LLM on user queries and GPT-4 generated responses across a wide range of real-world scenarios. It accommodates diverse evaluation protocols and has been validated in 58 different scenarios, even outperforming many proprietary LLMs. In light of increasingly complex user queries, we advocate that developing a more unified and contemporaneous evaluation protocol is a promising direction. Additionally, constructing high-quality, comprehensive datasets to train unified models holds great potential. Such advancements could significantly contribute to more effective and universal evaluations of LLMs.
@@ 6 Conclusion
In this survey, we have meticulously surveyed the role of LLMs in the evaluation of NLG. Our comprehensive taxonomy classifies works along three primary dimensions: evaluation function, evaluation references and evaluation task. This framework enabled us to systematically categorize and understand LLM-based evaluation methodologies. We delved into various LLM-based approaches, scrutinizing their strengths and comparing their differences. Additionally, we summarized prevalent meta-evaluation benchmarks for NLG evaluation. Throughout our study, we highlighted both the advancements and the prevailing challenges in this rapidly evolving field. While LLMs offer groundbreaking potential in evaluating NLG outputs, there still remain unresolved issues that require attention, including bias, robustness, the integration of hybrid evaluation methods, and the need for domain-specific and unified evaluation within LLM-based evaluators. We anticipate that addressing these challenges will pave the way for more general, effective, and reliable NLG evaluation techniques. Such advancements would contribute significantly to the progression of both NLG evaluation and the broader application of LLMs.

### v2

@@ 4 Benchmarks and Tasks
5 Comparison with Traditional Evaluators
Metrics Sup SummEval Topical-Chat WMT22
COH CON FLU REL Avg NAT COH ENG GRO Avg En-De En-Ru Zh-Eu
Traditional Metrics (Word Overlap)
ROUGE-1 0.167 0.160 0.115 0.326 0.192 0.158 0.206 0.319 0.264 0.233 - - -
ROUGE-2 0.184 0.187 0.159 0.290 0.205 0.168 0.247 0.337 0.311 0.266 - - -
ROUGE-L 0.128 0.115 0.105 0.311 0.165 0.145 0.205 0.306 0.293 0.237 - - -
BLEU - - - - - 0.175 0.235 0.316 0.310 0.259 0.169 0.140 0.145
BERT-based Metrics
BERTScore 0.284 0.110 0.193 0.312 0.225 0.209 0.233 0.335 0.317 0.273 0.232 0.192 0.316
BLEURT - - - - - - - - - - 0.344 0.359 0.361
BARTScore 0.448 0.382 0.356 0.356 0.385 -0.053 -0.079 -0.084 -0.197 -0.103 - - 0.220
UniTE - - - - - - - - - - 0.369 0.378 0.357
UniEval 0.575 0.446 0.449 0.426 0.474 0.450 0.616 0.615 0.590 0.568 - - -
LLM-based Metrics
GPTScore 0.434 0.449 0.403 0.381 0.417 - - - - - - - 0.187
CHATGPT(DA) 0.451 0.432 0.380 0.439 0.425 0.474 0.527 0.599 0.576 0.544 0.306 0.332 0.371
G-Eval 0.582 0.507 0.455 0.547 0.514 0.607 0.590 0.605 0.536 0.590 - - -
Embed Llama - - - - - - - - - - 0.400 0.227 0.217
X-Eval 0.530 0.428 0.461 0.500 0.480 0.478 0.622 0.593 0.728 0.605 - - -
Table 3: Performance of traditional and LLM-based metrics on Text Summarizing (SummEval), Dialogue Generation (Topical-Chat) and Machine Translation (WMT22) tasks. We demonstrate the sample-level Spearman correlations on SummEval and Topical-Chat benchmarks and the segment-level Kendall-Tau correlations on WMT22 benchmarks respectively. Sup indicates the metric is supervised. The specific representation of the evaluation aspects (COH/CON/FLU/REL/NAT/ENG/GRO) is shown in Table  3.1.
Qualitative Comparison
Traditional evaluation metrics (e.g., BLEU Papineni et al. (2002) and ROUGE) focus on exacting n-gram matches, which penalizes semantically correct but lexically different hypotheses. These methods are simple and fast but not robust to paraphrasing. BERTScore Zhang et al. (2020) measures quality through semantic similarity based on BERT contextual embeddings, effectively handling paraphrases and synonyms. However, such matching-based evaluators depend on the quality of the pre-trained embeddings, may struggle with very fine-grained semantic distinctions, and neglect the overall semantics of the hypotheses and references. Additionally, neither metric accounts for fluency or readability during evaluation and both still rely on reference texts.
In contrast, LLMs have a strong capability for language understanding and generation, which supports evaluating quality without needing references. They can adapt to various domains and languages, making them suitable for a wide range of NLG tasks without requiring task-specific feature engineering. LLMs also provide more nuanced evaluation criteria beyond traditional metrics, such as semantic coherence, fluency and possible explanations. However, LLM-based methods are computationally more intensive due to their vast architectures. Additionally, prompting LLMs for NLG evaluation requires careful crafting of prompts. Variations in these prompts can lead to substantial differences in evaluation outcomes, as indicated in Gao et al. (2023). Section 6 summarizes more open problems of LLM-based metrics.
Performance Comparison
Table 3 summarizes the performance of both traditional word-overlap metrics, BERT-based metrics and recent LLM-based metrics on representative benchmarks such as SummEval, WMT, and Topical-Chat. We can easy to observe that the latter two metrics generally perform better than word-overlap metrics. Despite not being fine-tuned, the most competitive LLM-based methods (e.g., G-Eval for summarization and CHATGPT(DA) for machine translation) generally achieve a higher correlation with all traditional metrics, whether for unsupervised or fine-tuned methods. These results reveal the strong capability of LLMs in language understanding, contextual analysis, coherence checking, and fluency assessment of generated text. Among the three tasks, the performance gap between LLM-based evaluators and traditional evaluators is not significant in the machine translation task. This phenomenon might be due to the limitations of LLM-based models in cross-lingual understanding. Additionally, according to the results of last row in the table, we can observe that the performance of different LLM-based metrics varies significantly, which implies their sensitivity to prompt crafting. In contrast, traditional unsupervised methods like ROUGE, BLEU, and BERTScore are more robust, although their overall performance is relatively worse.
Efficiency Comparison
Table 4 presents the average number of texts evaluated per second for different metrics in the SummEval (TS task) and Topical-chat (DG task) benchmarks. This comparison highlights the efficiency differences between traditional metrics and LLM-based metrics. Our tests were conducted on an NVIDIA A40 GPU. The results show that efficiency generally correlates with model size and traditional word-overlap metrics (e.g., BLEU and ROUGE) are significantly faster than other metrics. Specifically, LLM-based evaluators are about $200$ to $400$ times slower than traditional word-overlap metrics. However, their efficiency can be improved with advanced LM inference tools such as vLLM [^1] https://github.com/vllm-project/vllm. While LLM-based evaluators are suitable for offline evaluation, they may not be feasible for online evaluation.
Methods Backbone TS DG
BLEU - 977.31 2344.16
ROUGE - 446.36 2379.24
BERTScore BERT 37.64 42.37
ChatGPT(DA) ChatGPT 1.94 1.87
G-Eval GPT-4 1.51 1.40
TIGERScore Llama 2.67 3.72
Table 4: Efficiency Comparison. We report the aver- age number of texts evaluated per second for different metrics.
6 Challenges and Open Problems
@@
LLMs have been expanded w.r.t their broad capabilities beyond traditional single-task focuses, encompassing complex instructions like coding and open-ended real-world requirements OpenAI (2023); Significant Gravitas . Consequently, there is an increasing demand for more comprehensive and flexible evaluation methods. However, traditional evaluation methods and most current LLM-based evaluators are limited to constrained tasks and evaluation aspects (cf. Table 3.1). Some promising attempts have been made in this direction. For instance, MT-Bench Zheng et al. (2023) uses GPT-4 as an evaluator across multiple domains for multi-turn questions. Yet, this is too confined to a few evaluation aspects and limits dialogue to two turns only. Another model, Auto-J Li et al. (2023b), approaches from a data construction perspective, training a 13B LLM on user queries and GPT-4 generated responses across a wide range of real-world scenarios. It accommodates diverse evaluation protocols and has been validated in 58 different scenarios, even outperforming many proprietary LLMs. In light of increasingly complex user queries, we advocate that developing a more unified and contemporaneous evaluation protocol is a promising direction. Additionally, constructing high-quality, comprehensive datasets to train unified models holds great potential. Such advancements could significantly contribute to more effective and universal evaluations of LLMs.
@@ 7 Conclusion
In this paper, we have meticulously surveyed the role of LLMs in the evaluation of NLG. Our comprehensive taxonomy classifies works along three primary dimensions: evaluation function, evaluation references and evaluation task. This framework enabled us to systematically categorize and understand LLM-based evaluation methodologies. We delved into various LLM-based approaches, scrutinizing their strengths and comparing their differences. Additionally, we summarized prevalent meta-evaluation benchmarks for NLG evaluation. Throughout our study, we highlighted both the advancements and the prevailing challenges in this rapidly evolving field. While LLMs offer groundbreaking potential in evaluating NLG outputs, there still remain unresolved issues that require attention, including bias, robustness, the integration of hybrid evaluation methods, and the need for domain-specific and unified evaluation within LLM-based evaluators. We anticipate that addressing these challenges will pave the way for more general, effective, and reliable NLG evaluation techniques. Such advancements would contribute significantly to the progression of both NLG evaluation and the broader application of LLMs.